40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 22 02:21:57 2012, Jenne, Update, Green Locking, 2 arm ALS - Success!!!! P6210140.AVIspeedy_gonzales-5257.jpg
    Reply  Fri Jun 22 03:37:42 2012, Koji, Update, Green Locking, 2 arm ALS - Success!!!! 
       Reply  Fri Jun 22 10:52:18 2012, yuta, Update, Green Locking, 2 arm ALS - Success!!!! 
Message ID: 6851     Entry time: Fri Jun 22 02:21:57 2012     Reply to this: 6852
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: 2 arm ALS - Success!!!! 

[Yuta, Jenne]

We locked both arms using the ALS system simultaneously!  Hooray!

Video of spectrum analyzer during lock acquisition of both beats is attached.

Jamie is super awesome, since he fixed us up a beatbox speedy-quick.  Thanks Jamie!!  speedy_gonzales-5257.jpg

 Details:

1:  Aligned PSL green optics

     1.1:  We added an amplifier of ~20dB after the X beat PD (more Xgreen power on the PSL table so the signal was ~3dB higher than Y, so required less amplification).  The ~24dB amplifier is still in place after the Y beat PD.  Both beat signals go to a splitter after their amplifiers.  One side of each splitter goes to one of the channels on the beatbox.  The other side of each splitter goes to a 3rd splitter, which we're using backwards to combine the 2 signals so we can see both peaks on the spectrum analyzer at the same time.

2:  Found both beat notes

     2.1:  Y beat was easy since we knew the temps that have been working for the past several days

     2.2:  X beat was more tricky - the last time it was locked was the end of February (elog 6342)

         2.2.1:  We found it by adjusting the PSL laser temp nearly the full range - DC Adjust slider was at 8.8V or so (Y beat was found with the slider at ~1.1V tonight)

          2.2.2:  We then walked the beat around to get the PSL temp back to "normal" by moving the PSL temp, then compensating with the Xend laser temp, keeping the beatnote within the range of the spectrum analyzer.

          2.2.3:  Fine tuned the temps of all 3 lasers until we had 2 peaks on the analyzer at the same time!!

               2.2.3.1:  Yend - measured Temp=34.14 C, thermal Out of Slow servo=29820

               2.2.3.2:  Xend - displayed temp=39.33 C, thermal Out of Slow servo=5070

               2.2.3.3: PSL - displayed temp=31.49 C, Slow actuator Adjust=1.100V

3:  Locked both arms using ALS!!

     3.1:  We were a little concerned that the Xarm wasn't locking.  We tried switching the cables on the beatbox so that we used the old channels for the Xarm, since the old channels had been working for Y.  Eventually we discovered that the input of the filter module for ETMX's POS-ALS input was OFF, so we weren't really sending any signals to ETMX.  We reverted the cabling to how it was this evening when Jamie reinstalled the beatbox.

          3.1.1:  We need to sort out our SUS screens - Not all buttons in medm-land link to the same versions of the SUS screens!  It looks like the ALS screen was modified to point the ETMY button to a custom ETMY SUS screen which has the ALS path in the POS screen, along with LSC and SUSPOS.  There is no such screen (that I have found) for ETMX.  The regular IFO_ALIGN screen points to the generic SUS screens for both ETMY and ETMX, so we didn't know until Yuta searched around for the filter bank that the ALS input for ETMX was off.  We just need to make sure that all of the screens reflect what's going on in the models.

     3.2:  See the video attached - it shows the beat peaks during locking!!! (how do I embed it? right now you have to download it)

          3.2.1:  First you will see both peaks moving around freely

          3.2.2:  Then X arm is locked briefly, then unlocked

          3.2.3: Y arm is locked, steadily increasing gain

          3.2.4:  X arm is locked, so both arms locked simultaneously

          3.2.5:  Yuta clicked a button, accidentally unlocking the Xarm

4:  The transmission of the X arm was not so great, and both of our green beams (although X green especially) were no longer nicely aligned with the cavities.  Yuta tried to align the X arm to the X green, but it's bad enough that we really need to start over with the whole IFO alignment - we leave this until tomorrow.  Since we didn't have any good IR transmission, we didn't bother to try to find and hold the Xarm on IR resonance using ALS, so we didn't measure a POX out of loop residual cavity motion spectrum.  Again, tomorrow. 

Attachment 1: P6210140.AVI  22.271 MB
ELOG V3.1.3-